# Full-sky Models of Galactic Microwave Emission and Polarization at Sub-arcminute Scales for the Python Sky Model

### v1

@@ 2.2.5 CO Emission
In all CO models presented in this work, emission from the $J=1\rightarrow 0,2\rightarrow 1$, and $3\rightarrow 2$ transitions at 115.3, 230.5, and 345.8 GHz, respectively, is modeled as a delta function at the rest frequency. Therefore, a model is fully defined by maps of $I$, $Q$, and $U$ at the reference frequency. We present the new PySM CO models co1, co2, and co3 in Section 4.2.
2.2.6 Other Emission Mechanisms
@@
Other isotopologues of CO emit at frequencies near the 12C18O lines. However, these species produce much weaker emission and reside in even denser gas than does 12C18O, and so have a more minor role as a CMB foreground. Likewise, HCN has been identified as a potential contributor to the Commander sky model, but mostly toward the Galactic Center (Planck Collaboration X, 2016). Line emission from C II at 158 $\mu$m and N II at 122 and 205 $\mu$m from the diffuse ISM was mapped by COBE/FIRAS (Bennett et al., 1994), but there are currently no CMB experiments operating at such high frequencies. COBE/FIRAS also detected weaker line emission from C I (370 and 609 $\mu$m) and CO transitions beyond those modeled in this work, though these lines were mostly seen toward the Galactic Center (Bennett et al., 1994).
@@
3 Dust and Synchrotron Models: Stochastic Emission at Small Angular Scales
@@
The methods presented here aim to preserve the well-measured large-scale information in existing observations of dust and synchrotron emission (the “template” maps), filter out the noisy small-scale emission in the maps, and replace those small scales with a stochastic realization that has a reasonable correspondence with the large-scale emission. Specifically, the synthetic small-scale structure should have a power spectrum that connects smoothly to the power spectrum of the real data at large scales. Our approach is to generate stochastic realizations of small-scale emission that is modulated by the large-scale signal.
@@
We divide our discussion into amplitude parameters ($A_{d}^{S}$ and $A_{s}^{S}$ in Equations 1 and 2, respectively) in Section 3.1 and spectral parameters ($\beta_{d}$ and $T_{d}$ in Equation 1 and $\beta_{s}$ and $c_{s}$ in Equation 2) in Section 3.2 as the methodology differs for each. With maps of both sets of parameters, foreground model maps at any desired frequency can be generated from the amplitude maps at the reference frequency using Equations 1 and 2.
3.1 Small-Scale Fluctuations in Amplitude Parameters
@@
To implement the parametric models of dust and synchrotron emission described in Section 2.2, we require maps of $A^{I}$, $A^{Q}$, and $A^{U}$ for each mechanism, i.e., the total emission and polarization at a reference frequency. We start from template $I$, $Q$, and $U$ maps derived from data, to which we add fluctuations at scales that are noise-dominated.
3.1.1 The Polarization Fraction Tensor Formalism
@@
A principal challenge for generating realizations of Galactic emission at small angular scales over the full sky is that the amplitude of the fluctuations is strongly dependent on proximity to the Galactic plane. Further, Gaussian random fluctuations are a poor representation of the typically filamentary structure of the Galactic ISM (e.g., Hacar et al., 2023). We address each of these challenges through use of a polarization fraction tensor framework.
@@
While the mathematical model of this framework can be applied to any emission mechanism, it is motivated by a simple model of dust emission. In this picture, the morphology of the dust polarized intensity $P$ at large angular scales is set primarily by the density structure of the dust, probed in projection by the total intensity $I$, and then secondarily by the large-scale morphology of the Galactic magnetic field, which modulates the polarization fraction of the dust, $p\equiv P/I$. Random fluctuations of the turbulent component of the Galactic magnetic field lead to fluctuations on top of this smooth, large-scale polarized intensity distribution. The amplitude of these fluctuations is much more spatially isotropic than fluctuations in the $P$ map itself. Further, total and polarized dust emission are coupled through the angle $\gamma$ between the local magnetic field and the line of sight, such that the sum $I+P$ is independent of $\gamma$ and proportional to the dust column density (Hensley et al., 2019). Finally, Galactic dust emission has a large dynamic range in $I$, whereas $\ln I$ not only varies much less but is better described by a Gaussian distribution over the sky.
@@
A two-dimensional sky map of intensity and linear polarization is described by the symmetric rank-2 tensor (e.g., Landau & Lifshitz, 1975)
@@
where $\xi\equiv\sqrt{q^{2}+u^{2}}$. A number of useful features of this transformation are evident. First, $I$ is guaranteed to be positive for any values of $i$, $q$, and $u$. Further,
@@
3.1.2 General Methodology
Figure 1: An overview of the steps for building model d10. 1. We begin with GNILC-derived $I$, $Q$, $U$ maps of the polarized dust emission at 353 GHz (“amplitude parameters”, purple) and maps of the GNILC-derived dust temperature and spectral index (“spectral parameters”, brown). 2. The amplitude parameters are transformed via the polarization tensor fraction formalism to $i$, $q$, $u$. 3. Both sets of maps are filtered by scale: large scales are untouched, and small-scale map structure is replaced by structure synthesized from a power spectrum extrapolated from a power-law fit to the large-scale data. 4. The small-scale structure maps are modulated by maps of the large-scale structure in either total intensity or polarization: the fluctuations in $i$ and the spectral parameters are modulated by $m_{i}$, and the fluctuations in $q$ and $u$ are modulated by $m_{p}$. 5. The large- and small-scale maps are combined. 6. The $i$, $q$, $u$ maps are transformed back to $I$, $Q$, and $U$. 7. Together, the final amplitude maps at 353 GHz, and the final spectral parameter maps, which define how the sky varies with frequency, fully specify the d10 model.
@@ 3.1.2 General Methodology
    We first transform the $I$, $Q$, and $U$ templates into $i$, $q$, and $u$ templates via Equation (3.1.1).
@@
    We then compute the $tt$, $te$, $ee$, and $bb$ full-sky power spectra from the $i$, $q$, and $u$ maps in analogy with how $TT$, $TE$, $EE$, and $BB$ spectra are computed from $I$, $Q$, and $U$ maps.
@@
    We synthesize $i$, $q$, and $u$ maps using the constructed $tt$, $te$, $ee$, and $bb$ spectra. These maps are then high-pass filtered at cut-off multipole $\ell_{1}$.
@@
    Finally, we combine the large-scale and small-scale maps (from Steps 2 and 6, respectively) and transform back to $I$, $Q$, and $U$ maps via Equation (3.1.1).
@@
To construct the modulation maps $m_{i}$ and $m_{p}$, we generally follow the method outlined in Thorne et al. (2017). For each pixel at the location $\hat{\theta}$ in an $N_{\rm side}=8$ HEALPix map, we define a circular mask centered on $\hat{\theta}$ of radius $11.3^{\circ}$ and apodize it with a $2^{\circ}$ Gaussian profile. We compute the $tt$ and $ee$ power spectra within each circle using NaMaster (Alonso et al., 2019), to which we fit simple power laws in $\ell$. We then evaluate these best-fit power-law spectra at the $\ell_{*}$ multipole (Table 1). Finally, the modulation maps are constructed from the ratios:
$m_{i}\left(\hat{\theta}\right)$ $=\left(\frac{C^{tt}_{\ell_{*},circ}}{C^{tt}_{\ell_{*},full}}\right)^{1/2}$ (8)
$m_{p}\left(\hat{\theta}\right)$ $=\left(\frac{C^{ee}_{\ell_{*},circ}}{C^{ee}_{\ell_{*},full}}\right)^{1/2}$ (9)
@@
Whereas previous PySM models used modulation maps constructed in $N_{\rm side}=2$ superpixels and then smoothed with a $10^{\circ}$ beam (Thorne et al., 2017), we employ $N_{\rm side}=8$ superpixels and overlapping circular regions. This is motivated by improvements to the Planck polarization maps since the earlier study. Thus the models developed in this work incorporate smaller-scale information into the modulation maps than previous PySM models.
The adopted values of $\ell_{1}$ and $\ell_{*}$ are driven both by the angular resolution and the signal-to-noise ratio of the template maps. We employ $\ell_{1}=100$ and $\ell_{*}=80$ for dust and $\ell_{1}=38$ and $\ell_{*}=36$ for synchrotron. The $\ell_{2}$ parameter is chosen to be high enough that, for $\ell>\ell_{2}$, dust and synchrotron emission are relatively unconstrained by current full-sky data. We adopt $\ell_{2}=2000$ for dust and $\ell_{2}=400$ for synchrotron. For dust we employ a power law index of $\alpha_{tt}=-0.80$ based on the range of fit $\alpha_{TT}$ in Planck, WISE, and MegaCam data by Miville-Deschênes et al. (2016). We adopt $\alpha_{ee}=-0.42$, $\alpha_{te}=-0.50$, and $\alpha_{bb}=-0.54$ from fits to Planck $TT$, $TE$, and $BB$ spectra, respectively, over 71% of the sky (Planck Collaboration XI, 2020). For synchrotron we adopt $\alpha_{tt}=-1.00$ and $\alpha_{te}=-1.00$ based on the observed spectra of the synchrotron template map. We adopt $\alpha_{ee}=-0.84$ and $\alpha_{bb}=-0.76$ following fits to synchrotron maps over 78% of the sky produced with the Commander pipeline (Planck Collaboration IV, 2020). The $\ell_{1}$, $\ell_{2}$, $\ell_{*}$, and $\alpha$ values used to construct the small-scale maps are all summarized in Table 1.
@@
Another challenge for our prescription is pixels that, whether from noise or systematic errors, have $I^{2}<P^{2}$, rendering the transformation in Equation 3.1.1 invalid. We identify 527 such pixels in our dust template and 5 in our synchrotron template, all in the vicinity of the Crab Nebula. To address this, we set the $i$, $q$, and $u$ values at the location of these pixels to fiducial values of 4.51, 0.07, and 0.01 in the dust maps and 4.55, 0.017, $-0.001$ in the synchrotron maps. Given these complications, we do not advise using the simulations in the vicinity of the Crab Nebula.
Finally, we emphasize that our procedures are optimized for producing polarization rather than total intensity maps. For instance, we could in principle retain much smaller angular scales from the intensity templates, however we require that the $\ell_{0}$, $\ell_{1}$, $\ell_{2}$, and $\ell_{*}$ parameters are the same in total intensity and polarization in order to implement the polarization fraction tensor framework. Other choices, such as the value of $\gamma$, are based on polarization spectra. Some of the imperfections of the maps in total intensity produced as a consequence of these choices are documented in Section 5.
@@ 3.1.3 Dust Amplitude
The Planck 2015 component separation results in total intensity remain state of the art despite updates in polarization in the 2018 data release (Planck Collaboration IV, 2020). Previous PySM models, e.g., d0 and d1, employed the dust templates from the Commander component separation analysis (Planck Collaboration IX, 2016). However, the model fitting employed in Planck Collaboration IX (2016) did not differentiate between Galactic dust emission and the Cosmic Infrared Background (CIB), and so the component separated dust maps retain CIB signal that should not be included in simulations of Galactic emission (see Section 5.4 for a detailed discussion). To address this issue, we instead use dust templates from analyses that separated Galactic dust emission from the CIB using the Generalized Needlet Internal Linear Combination (GNILC) algorithm (Remazeilles et al., 2011).
@@
In total intensity we employ the Planck GNILC 2015 component separated map [^6] COM_CompMap_Dust-GNILC-F353_2048_R2.00.fits at $353$ GHz (Planck Collaboration Int. XLVIII, 2016). This map has variable angular resolution, ranging from $21.8\arcmin$ to $5\arcmin$ depending on the sky region, which complicates its use in our analysis. Therefore, we reprocess the map to a spatially uniform resolution of $21.8\arcmin$. To do so, we first note that the map was synthesized from ten needlet maps of different, but spatially uniform, resolutions (Planck Collaboration Int. XLVIII, 2016, Figure A.2). Each of the $j=1,\cdots,10$ needlet maps can be reconstructed from the Planck GNILC 2015 dust map as follows:
$\gamma^{j}(\hat{n})=\sum_{\begin{subarray}{c}\ell,m\end{subarray}}\left(a_{\ell m}\,h^{j}_{\ell}\right)Y_{\ell m}(\hat{n})\,,$ (12)
where $a_{\ell m}$ are the spherical harmonic coefficients of the Planck GNILC 2015 dust map, $h^{j}_{\ell}$ is the $j^{\text{th}}$ needlet window function (see Planck Collaboration Int. XLVIII, 2016, Figure A.2), and $Y_{\ell m}(\hat{n})$ denotes spherical harmonics. We produce a map of uniform resolution by retaining only the first six needlet maps [^7] Available at NERSC: https://portal.nersc.gov/project/cmb/pysm-data/dust_gnilc/inputs/, which contain full-sky information and probe the dust intensity from the largest scales down to $21.8\arcmin$:
$d^{\text{GNILC,}\,21.8^{\prime}}(\hat{n})=\sum_{j=1}^{6}h^{j}_{\ell}\gamma^{j}(\hat{n})\,.$ (13)
@@
This yields an $N_{\rm side}=2048$ map that reproduces the Planck GNILC 2015 dust intensity template at all scales $\geq 21.8\arcmin$. Finally, we subtract the CIB monopole of $0.13\,\text{MJy}\,\text{sr}^{-1}$ present in the map (Planck Collaboration XII, 2020, Section 2.2).
@@
For the dust $Q$ and $U$ maps we employ the dust maps [^8] COM_CompMap_IQU-thermaldust-gnilc-varres_2048_R3.00.fits produced by the GNILC component separation from the Planck Public Release 3 (Planck Collaboration IV, 2020; Planck Collaboration XII, 2020). While these maps have variable angular resolution, the resolution varies over the sky much more smoothly than in the $I$ map. Given this, and that we wish to retain as much polarization information in our analysis as possible, we employ these variable resolution maps as-is. Their resolution ranges from $5\arcmin$ in the Galactic plane to $80\arcmin$ at high Galactic latitudes. The maps are pixelated at $N_{\rm side}=2048$.
@@
To produce dust emission templates at a monochromatic frequency of 353 GHz, we divide each of the $I$, $Q$, and $U$ maps by a factor of 1.098 to correct for the Planck bandpass (Planck Collaboration XI, 2020, Table 2).
3.1.4 Synchrotron Amplitude

### v2

@@ II.2.5 CO Emission
In all CO models presented in this work, emission from the $J=1\rightarrow 0,2\rightarrow 1$, and $3\rightarrow 2$ transitions at 115.3, 230.5, and 345.8 GHz, respectively, is modeled as a delta function at the rest frequency. Therefore, a model is fully defined by maps of $I$, $Q$, and $U$ at the reference frequency. We present the new PySM CO models co1, co2, and co3 in Section IV.2.
II.2.6 Other Emission Mechanisms
@@
[…]
@@
III Dust and Synchrotron Models: Stochastic Emission at Small Angular Scales
@@
The methods presented here aim to preserve the well-measured large-scale information in existing observations of dust and synchrotron emission (the “template” maps), filter out the noisy small-scale emission in the maps, and replace those small scales with a stochastic realization that has a reasonable correspondence with the large-scale emission. Specifically, the synthetic small-scale structure should have a power spectrum that connects smoothly to the power spectrum of the real data at large scales. We rely on the fact that the angular power spectra of both synchrotron and thermal dust emission are well-approximated by a power-law in $\ell$ (planck2014-XIX; Miville-Deschenes:2016). Our approach is to generate stochastic realizations of small-scale emission that is modulated by the large-scale signal.
@@
We divide our discussion into amplitude parameters ($A_{d}^{S}$ and $A_{s}^{S}$ in Equations 1 and 2, respectively) in Section III.1 and spectral parameters ($\beta_{d}$ and $T_{d}$ in Equation 1 and $\beta_{s}$ and $c_{s}$ in Equation 2) in Section III.2 as the methodology differs for each. With maps of both sets of parameters, foreground model maps at any desired frequency can be generated from the amplitude maps at the reference frequency using Equations 1 and 2.
III.1 Small-Scale Fluctuations in Amplitude Parameters
@@
To implement the parametric models of dust and synchrotron emission described in Section II.2, we require maps of $A^{I}$, $A^{Q}$, and $A^{U}$ for each mechanism, i.e., the total emission and polarization at a reference frequency. We start from template $I$, $Q$, and $U$ maps derived from data, to which we add fluctuations at scales that are noise-dominated.
III.1.1 The Polarization Fraction Tensor Formalism
@@
A principal challenge for generating realizations of Galactic emission at small angular scales over the full sky is that the amplitude of the fluctuations is strongly dependent on proximity to the Galactic plane. Further, Gaussian random fluctuations are a poor representation of the typically filamentary structure of the Galactic ISM (e.g., Hacar:2023). We address each of these challenges through use of a polarization fraction tensor framework.
@@
While the mathematical model of this framework can be applied to any emission mechanism, it is motivated by a simple model of dust emission. In this picture, the morphology of the dust polarized intensity $P$ at large angular scales is set primarily by the density structure of the dust, probed in projection by the total intensity $I$, and then secondarily by the large-scale morphology of the Galactic magnetic field, which modulates the polarization fraction of the dust, $p\equiv P/I$. Random fluctuations of the turbulent component of the Galactic magnetic field lead to fluctuations on top of this smooth, large-scale polarized intensity distribution. The amplitude of these fluctuations is much more spatially isotropic than fluctuations in the $P$ map itself. Further, total and polarized dust emission are coupled through the angle $\gamma$ between the local magnetic field and the line of sight, such that the sum $I+P$ is independent of $\gamma$ and proportional to the dust column density (Hensley:2019). Finally, Galactic dust emission has a large dynamic range in $I$, whereas $\ln I$ not only varies much less but is better described by a Gaussian distribution over the sky.
@@
A two-dimensional sky map of intensity and linear polarization is described by the symmetric rank-2 tensor (e.g., Landau:1975)
@@
where $\xi\equiv\sqrt{q^{2}+u^{2}}$. A number of useful features of this transformation are evident. First, II is guaranteed to be positive for any values of $i$, $q$, and $u$. Further,
@@
III.1.2 General Methodology
Figure 1: An overview of the steps for building model d10. 1. We begin with GNILC-derived II, $Q$, $U$ maps of the polarized dust emission at 353 GHz (“amplitude parameters”, purple) and maps of the GNILC-derived dust temperature and spectral index (“spectral parameters”, brown). 2. The amplitude parameters are transformed via the polarization tensor fraction formalism to $i$, $q$, $u$. 3. Both sets of maps are filtered by scale: large scales are untouched, and small-scale map structure is replaced by structure synthesized from a power spectrum extrapolated from a power-law fit to the large-scale data. 4. The small-scale structure maps are modulated by maps of the large-scale structure in either total intensity or polarization: the fluctuations in $i$ and the spectral parameters are modulated by $m_{i}$, and the fluctuations in $q$ and $u$ are modulated by $m_{p}$. 5. The large- and small-scale maps are combined. 6. The $i$, $q$, $u$ maps are transformed back to $I$, $Q$, and $U$. 7. Together, the final amplitude maps at 353 GHz, and the final spectral parameter maps, which define how the sky varies with frequency, fully specify the d10 model.
@@ III.1.2 General Methodology
    We first transform the $I$, $Q$, and $U$ templates into $i$, $q$, and $u$ templates via Equation (III.1.1).
@@
    We then compute the $tt$, tete, eeee, and $bb$ full-sky power spectra from the $i$, $q$, and $u$ maps in analogy with how $TT$, TETE, EEEE, and $BB$ spectra are computed from $I$, $Q$, and $U$ maps.
@@
    We synthesize $i$, $q$, and $u$ maps using the constructed $tt$, tete, eeee, and $bb$ spectra. These maps are then high-pass filtered at cut-off multipole $\ell_{1}$.
@@
    Finally, we combine the large-scale and small-scale maps (from Steps 2 and 6, respectively) and transform back to $I$, $Q$, and $U$ maps via Equation (III.1.1).
@@
To construct the modulation maps $m_{i}$ and $m_{p}$, we generally follow the method outlined in Thorne:2017. For each pixel at the location $\hat{\theta}$ in an $N_{\rm side}=8$ HEALPix map, we define a circular mask centered on $\hat{\theta}$ of radius $11.3^{\circ}$ and apodize it with a $2^{\circ}$ Gaussian profile. We compute the $tt$ and $ee$ power spectra within each circle using NaMaster (Alonso:2019), to which we fit simple power laws in $\ell$. We then evaluate these best-fit power-law spectra at the $\ell_{*}$ multipole (Table 1). Finally, the modulation maps are constructed from the ratios:
[…]
@@
Whereas previous PySM models used modulation maps constructed in $N_{\rm side}=2$ superpixels and then smoothed with a $10^{\circ}$ beam (Thorne:2017), we employ $N_{\rm side}=8$ superpixels and overlapping circular regions. This is motivated by improvements to the Planck polarization maps since the earlier study. Thus the models developed in this work incorporate more small-scale information into the modulation maps than previous PySM models.
The adopted values of $\ell_{1}$ and $\ell_{*}$ are driven both by the angular resolution and the signal-to-noise ratio of the template maps. We employ $\ell_{1}=100$ and $\ell_{*}=80$ for dust and $\ell_{1}=38$ and $\ell_{*}=36$ for synchrotron. The $\ell_{2}$ parameter is chosen to be high enough that, for $\ell>\ell_{2}$, dust and synchrotron emission are relatively unconstrained by current full-sky data. We adopt $\ell_{2}=2000$ for dust and $\ell_{2}=400$ for synchrotron. For dust we employ a power law index of $\alpha_{tt}=-0.80$ based on the range of fit $\alpha_{TT}$ in Planck, WISE, and MegaCam data by Miville-Deschenes:2016. We adopt $\alpha_{ee}=-0.42$, $\alpha_{te}=-0.50$, and $\alpha_{bb}=-0.54$ from fits to Planck $TT$, TETE, and $BB$ spectra, respectively, over 71% of the sky (planck2016-l11A). For synchrotron we adopt $\alpha_{tt}=-1.00$ and $\alpha_{te}=-1.00$ based on the observed spectra of the synchrotron template map. We adopt $\alpha_{ee}=-0.84$ and $\alpha_{bb}=-0.76$ following fits to synchrotron maps over 78% of the sky produced with the Commander pipeline (planck2016-l04). The $\ell_{1}$, $\ell_{2}$, $\ell_{*}$, and $\alpha$ values used to construct the small-scale maps are all summarized in Table 1.
@@
Another challenge for our prescription is pixels that, whether from noise or systematic errors, have $I^{2}<P^{2}$, rendering the transformation in Equation III.1.1 invalid. We identify 527 such pixels in our dust template and 5 in our synchrotron template, all in the vicinity of the Crab Nebula. To address this, we set the $i$, $q$, and $u$ values at the location of these pixels to fiducial values of 4.51, 0.07, and 0.01 in the dust maps and 4.55, 0.017, 0.001-0.001 in the synchrotron maps. Given these complications, we do not advise using the simulations in the vicinity of the Crab Nebula.
Finally, we emphasize that our procedures are optimized for producing polarization rather than total intensity maps. For instance, we could in principle retain much smaller angular scales from the intensity templates, however we require that the $\ell_{0}$, $\ell_{1}$, $\ell_{2}$, and $\ell_{*}$ parameters are the same in total intensity and polarization in order to implement the polarization fraction tensor framework. Other choices, such as the value of $\gamma$, are based on polarization spectra. Some of the imperfections of the maps in total intensity produced as a consequence of these choices are documented in Section V.
@@ III.1.3 Dust Amplitude
The Planck 2015 component separation results in total intensity remain state of the art despite updates in polarization in the 2018 data release (planck2016-l04). Previous PySM models, e.g., d0 and d1, employed the dust templates from the Commander component separation analysis (planck2014-a11). However, the model fitting employed in planck2014-a11 did not differentiate between Galactic dust emission and the Cosmic Infrared Background (CIB), and so the component separated dust maps retain CIB signal that should not be included in simulations of Galactic emission (see Section V.4 for a detailed discussion). To address this issue, we instead use dust templates from analyses that separated Galactic dust emission from the CIB using the Generalized Needlet Internal Linear Combination (GNILC) algorithm (Remazeilles:2011).
@@
In total intensity we employ the Planck GNILC 2015 component separated map [^6] COM_CompMap_Dust-GNILC-F353_2048_R2.00.fits at $353$ GHz (planck2016-XLVIII). This map has variable angular resolution, ranging from $21.8\arcmin$ to $5\arcmin$ depending on the sky region, which complicates its use in our analysis. Therefore, we reprocess the map to a spatially uniform resolution of $21.8\arcmin$. To do so, we first note that the map was synthesized from ten needlet maps of different, but spatially uniform, resolutions (planck2016-XLVIII, Figure A.2). Each of the $j=1,\cdots,10$ needlet maps can be reconstructed from the Planck GNILC 2015 dust map as follows:
[…]
@@
This yields an $N_{\rm side}=2048$ map that reproduces the Planck GNILC 2015 dust intensity template at all scales $\geq 21.8\arcmin$. Finally, we subtract the CIB monopole of $0.13\,\text{MJy}\,\text{sr}^{-1}$ present in the map (planck2016-l11B, Section 2.2).
@@
For the dust $Q$ and $U$ maps we employ the dust maps [^8] COM_CompMap_IQU-thermaldust-gnilc-varres_2048_R3.00.fits produced by the GNILC component separation from the Planck Public Release 3 (planck2016-l04; planck2016-l11B). While these maps have variable angular resolution, the resolution varies over the sky much more smoothly than in the $I$ map. Given this, and that we wish to retain as much polarization information in our analysis as possible, we employ these variable resolution maps as-is. Their resolution ranges from $5\arcmin$ in the Galactic plane to $80\arcmin$ at high Galactic latitudes. The maps are pixelated at $N_{\rm side}=2048$.
@@
To produce dust emission templates at a monochromatic frequency of 353 GHz, we divide each of the $I$, $Q$, and $U$ maps by a factor of 1.098 to correct for the Planck bandpass (planck2016-l11A, Table 2).
III.1.4 Synchrotron Amplitude